# EcoLoRA: Communication-Efficient Federated Fine-Tuning of Large Language Models

### v1

@@ Abstract
To address data locality and privacy restrictions, Federated Learning (FL) has recently been adopted to fine-tune large language models (LLMs), enabling improved performance on various downstream tasks without requiring aggregated data. However, the repeated exchange of model updates in FL can result in prohibitively high communication costs, hindering the distributed learning process. To address this challenge, we propose EcoLoRA, a novel communication-efficient federated fine-tuning framework for LLMs. Leveraging the modular structure, we propose a round-robin segment sharing scheme, where each client uploads only a complementary LoRA segment per round to reduce network bandwidth. It is further combined with adaptive sparsification methods tailored to LoRA’s training dynamics and lossless encoding techniques. We conduct extensive evaluations on both question-answering and value-alignment tasks across multiple datasets and models. The results show that EcoLoRA significantly reduces communication overhead without compromising performance. For instance, it reduces communication time by up to 79% and total training time by up to 65%.
@@ 1 Introduction
With the advancements in scaling laws Kaplan et al. (2020), the parameter sizes of pre-trained language models have grown exponentially Chowdhery et al. (2023). Despite this rapid expansion, large language models (LLMs) remain constrained by their inherent knowledge boundaries, limiting their effectiveness in certain downstream tasks Mao et al. (2024). These limitations necessitate task-specific fine-tuning. However, the substantial data required for fine-tuning is often distributed across multiple entities, raising significant privacy concerns related to data sharing.
@@
To demonstrate the effectiveness of EcoLoRA, we incorporate it into various state-of-the-art methods across different tasks (including general question answering and value alignment), datasets, and models. Our results show significant communication savings while preserving model performance. Notably, EcoLoRA reduces uploaded parameters by up to 89% and overall communication parameters by up to 58% compared to existing approaches. Under practical network conditions, it reduces communication time by up to 79% and total training time by 65%. Moreover, our approach remains robust under various non-i.i.d. data conditions and adds only minimal computational overhead.
@@ 2.1 Parameter-efficient Fine-tuning of LLMs
The large parameter size of LLMs often makes traditional full fine-tuning prohibitively expensive. To address this challenge, various parameter-efficient fine-tuning (PEFT) techniques have been proposed, including prefix-tuning Li and Liang (2021), prompt-tuning Lester et al. (2021), and adapter-based methods Hu et al. (2023). Among these approaches, low-rank adaptation (LoRA) Hu et al. (2022), which leverages low-rank matrices to re-parameterize pre-trained weight matrices, has received unprecedented attention. LoRA requires tuning less than 1% of the parameters needed for a full fine-tune while still achieving comparable performance across a wide range of downstream tasks, without introducing additional inference latency. Building on these advantages, numerous LoRA variants have been developed to further improve its efficiency and accuracy Kopiczko et al. (2023); Zhang et al. (2023); Liu et al. (2024).
@@ 2.2 Federated Fine-tuning of LLMs
Federated fine-tuning of LLMs has recently received significant attention. Most of the existing work has focused on integrating PEFT methods into federated learning to reduce computation costs Che et al. (2023); Wu et al. (2024); Cho et al. (2024); Babakniya et al. ; Zhang et al. (2024); Sun et al. (2024); Bai et al. (2024). For example, Zhang et al. (2024) incorporate LoRA into the FedAvg framework so that only LoRA modules need to be trained and aggregated. Extending this approach to resource-constrained and heterogeneous scenarios, Wang et al. (2024) propose a stacking-based aggregation strategy for heterogeneous LoRA modules, where individual LoRA modules are uploaded for aggregation, and the resulting stacked full-size LoRA weights are distributed back to clients. Sun et al. (2024) further enhance performance under differential privacy guarantees and improve computational efficiency by fine-tuning only the zero-initialized LoRA matrices. Although these approaches reduce both computation and communication costs compared to full fine-tuning, transmitting LoRA modules still imposes considerable overhead. Even though LoRA accounts for a small portion of the total parameters, the massive scale of LLMs means these modules remain large. Repeatedly exchanging them during multiple communication rounds results in prohibitively high communication costs, making communication the dominant bottleneck in training time.
Another line of research leverages zeroth-order optimization methods for federated LLM fine-tuning Qin et al. (2024); Xu et al. (2024). While these approaches improve communication efficiency, their reliance on zeroth-order optimization significantly reduces computational efficiency compared to backpropagation-based methods. Consequently, these techniques substantially increase the computation time and, therefore, prolong the overall training process, particularly in scenarios with limited participating clients or resource-constrained environments.
@@ 3 Method
3.1 Problem Formulation
We consider an FL setting with one server and $K$ devices. Each device $i$ holds a local dataset $\mathcal{D}_{i}={(x_{j},y_{j})}^{n_{i}}$, where $n_{i}$, $x_{j}$, $y_{j}$ denote the number of samples, the input samples, and labels in client $i$, respectively. The total number of samples across all devices is $N=\sum_{i=1}^{K}n_{i}$. Following recent state-of-the-art approaches, the pre-trained LLMs $\mathcal{M}$ remain fixed on each device, while only the LoRA parameters are updated and exchanged between device $i$ and the server. The optimization goal is to find a set of LoRA parameters $\mathcal{P}$ to minimize the global loss:
where $\mathcal{L}(\mathcal{M},\mathcal{P},x_{j},y_{j})$ is the loss evaluated by the model $\mathcal{M}$ with LoRA parameters $\mathcal{P}$ on the local data $(x_{j},y_{j})$. LoRA models the weight update $\Delta W\in\mathbb{R}^{m\times n}$ through a low-rank decomposition $BA$, where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are two low-rank matrices with $r\ll\min(m,n)$.
@@ 3.2 System Model
Our primary objective is to enhance communication efficiency specifically for federated fine-tuning of LLMs. This setting, however, differs significantly from traditional FL paradigms, particularly due to the use of parameter-efficient fine-tuning techniques, which result in distinct training dynamics and parameter distributions. As a result, conventional communication optimization methods, such as top-$k$ sparsification, may fail to fully leverage these unique properties, leading to suboptimal communication gains. Other methods, like active client sampling Tang et al. (2022), may introduce considerable computational overhead, which undermines their practicality in LLM fine-tuning scenarios. These limitations highlight the need for communication-efficient techniques specifically tailored to the characteristics of federated LLM fine-tuning. To guide the design of such methods, we establish the following system goals:
@@
  • Robustness to Non-i.i.d. Data: Because data distributions can vary significantly across clients in real-world settings, our framework should remain robust under non-i.i.d. conditions.
To address these challenges, we propose EcoLoRA, a novel communication-efficient FL framework illustrated in Figure 1. First, we propose a round-robin segment-sharing scheme, leveraging the modular structure of LoRA. Instead of transmitting the entire LoRA module, each client shares only a complementary portion, significantly reducing communication overhead. Second, we introduce an adaptive sparsification method tailored for the different training dynamics observed in matrices A and B. This method dynamically compresses parameters based on their specific training behavior, ensuring minimal performance degradation. Third, the adaptive sparsification method naturally enables a parameter distribution suitable for geometric compression, which we exploit through Golomb coding to further optimize communication efficiency. We elaborate round-robin segment sharing, adaptive sparsification, and encoding in Sections 3.3, 3.4, and 3.5, respectively. Additionally, we analyze computational overhead in Section 3.6 and provide convergence analysis in Section 3.7.
@@ 3.3 Round-Robin Segment Sharing
For example, consider $N_{t}=5$ clients and $N_{s}=3$ segments. In round $t=0$, client 00 uploads the segment with ID $(0+0)\bmod 3=0$, i.e., $s_{0,0}^{0}$; client $1$ uploads $s_{1,1}^{0}$; client $2$ uploads $s_{2,2}^{0}$; client $3$ uploads $s_{3,0}^{0}$; and client $4$ uploads $s_{4,1}^{0}$. The server then averages $s_{0,0}^{0}$ and $s_{3,0}^{0}$ to form the 0-th segment, averages $s_{1,1}^{0}$ and $s_{4,1}^{0}$ to form the 1-th segment, and takes $s_{2,2}^{0}$ for the 2-th segment. Because each client transmits only a single segment in each round, this round-robin segment sharing scheme reduces the upload communication load to $1/N_{s}$ of the total parameters.
However, this partial update approach introduces a delay for segments that are not uploaded in a given round, which can increase the number of rounds required to converge. To mitigate potential accuracy degradation, we leverage the local model by taking a weighted average of the global and local models at the beginning of each round before optimization. This ensures that even if a segment is not uploaded in a particular round, its previous state still guides local optimization. Moreover, by mixing the globally shared model (the consensus among clients) with the client’s locally fine-tuned model (adapted to its specific data), we improve robustness under non-i.i.d. distributions. In cross-device settings, only a subset of clients participates in each round, which may result in some clients remaining idle for many rounds and thus suffering from stale local parameters that potentially hamper global convergence Xie et al. (2019) when using the simple average. To address this, we employ an exponential decay weighting Chen et al. (2019) to update the local LoRA model:
@@ 3.4 Adaptive Sparsification
where $k^{t}$ is the sparsity level for round $t$, $L_{0}$ is the initial loss, $L_{t-1}$ is the global loss for round $t-1$, and $k_{\max}$ and $k_{\min}$ define the sparsification range. As training loss decreases, $k^{t}$ is reduced, reflecting that the model has learned sufficient knowledge and updates have become sparser. Second, to address the distinct patterns in matrices A and B, we introduce a matrix-adaptive sparsification scheme. We set smaller $k_{\min}$ value for B (due to its higher sparsity) and use a larger $\gamma$ for B to capture its rapid change in sparsity.
@@ 3.5 Lossless Encoding
To communicate the set of sparse LoRA tensors between the server and the client, we only need to transmit the positions of the nonzero elements in the flattened tensors, along with a one-bit sign and 16-bit values (assuming FP16) for each nonzero update. However, the positions can still be expensive to communicate because they are typically stored with a fixed number of 16 bits. From an information-theoretic perspective, we can further compress these positions using lossless encoding Sattler et al. (2019). Specifically, rather than sending the absolute positions of the nonzero elements, we send the distances between consecutive nonzero positions. Given our adaptive sparsification rate $k$, each element is nonzero with probability $k$, thus the distance between two consecutive nonzero elements follows a geometric distribution with parameter $k$, where the probability of a distance of length $n$ is $(1-k)^{n-1}k$. For such geometrically distributed nonnegative integers, Golomb coding provides an optimal encoding scheme by representing each number as a combination of quotient and remainder Golomb (1966). For example, when $k=0.1$, using Golomb coding can reduce the average number of bits required to encode each nonzero position to $b^{*}=4.8$, which leads to approximately a 3.3$\times$ compression factor per position.
@@ 4.1 Experimental Setup
FL Settings and Implementation Details. Following Zhang et al. (2024), we implement our framework in a federated learning environment with 100 clients. In each round, we randomly sample 10 clients and conduct training for 40 global rounds. To simulate realistic scenarios, we adopt a non-i.i.d. data distribution across clients. Detailed experimental configurations and hyperparameter settings are provided in Appendix A [^1] We will release our code upon publication of the paper..
@@ 4.3 Evaluation in Practical Networks
To evaluate the performance of EcoLoRA under realistic network conditions, we implemented a simulated federated learning platform following Ekaireb et al. (2022), using ns-3, a widely adopted discrete-event simulator for network communications Henderson et al. (2008).
Following practical uplink (UL) and downlink (DL) bandwidth settings in Konečnỳ (2016), we simulate four bandwidth scenarios: 0.2/1 Mbps, 1/5 Mbps, 2/10 Mbps, and 5/25 Mbps, with a fixed latency of 50ms to capture different network conditions. Figure 3 compares the computation and communication time of EcoLoRA against baselines under these scenarios, using Llama2-7B trained on Dolly. Our results demonstrate that as network conditions deteriorate, communication time increasingly dominates the total training time. This effect is particularly notable given that actual throughput typically falls short of theoretical bandwidth. These findings underscore the importance of developing communication-efficient fine-tuning methods. Across all conditions, EcoLoRA significantly reduces communication overhead while introducing minimal computational cost. For instance, under the 1/5 Mbps setting, it reduces communication time by 79% and total training time by 65%. Moreover, the additional per-round computation cost remains below 3s, making EcoLoRA a practical solution for resource-constrained environments.
@@ 4.4 Ablation Study
In this section, we analyze the impact of various design components and hyperparameter choices. We also provide a comprehensive comparison with top-$k$ sparsification and present additional experiments under non-i.i.d. conditions in Appendix C.

### v2

@@ Abstract
To address data locality and privacy restrictions, Federated Learning (FL) has recently been adopted to fine-tune large language models (LLMs), enabling improved performance on various downstream tasks without requiring aggregated data. However, the repeated exchange of model updates in FL can result in prohibitively high communication costs, hindering the distributed learning process.
To address this challenge, we propose EcoLoRA, a novel communication-efficient federated fine-tuning framework for LLMs. Leveraging the modular structure, we propose a round-robin segment sharing scheme, where each client uploads only a complementary LoRA segment per round to reduce network bandwidth. It is further combined with adaptive sparsification methods tailored to LoRA’s training dynamics and lossless encoding techniques. We conduct extensive evaluations on both question-answering and value-alignment tasks across multiple datasets and models. The results show that EcoLoRA significantly reduces communication overhead without compromising performance. For instance, it reduces communication time by up to 79% and total training time by up to 65%.
@@ 1 Introduction
With the advancements in scaling laws Kaplan et al. (2020), the parameter sizes of pre-trained language models have grown exponentially Chowdhery et al. (2023). Despite this rapid expansion, large language models (LLMs) remain constrained by their inherent knowledge boundaries, limiting their effectiveness in certain downstream tasks Liu et al. (2024a). These limitations necessitate task-specific fine-tuning. However, the substantial data required for fine-tuning is often distributed across multiple entities, raising significant privacy concerns related to data sharing.
@@
To demonstrate the effectiveness of EcoLoRA, we incorporate it into various state-of-the-art methods across different tasks (including general question answering and value alignment), datasets, and models. Our results show significant communication savings while preserving model performance. Notably, EcoLoRA reduces uploaded parameters by up to 89% and overall communication parameters by up to 58% compared to existing approaches. Under practical network conditions, it reduces communication time by up to 79% and total training time by 65%. Moreover, our approach remains robust under various non-IID data conditions and adds only minimal computational overhead.
@@ 2.1 Parameter-efficient Fine-tuning of LLMs
Fine-tuning is essential for effectively adapting LLMs to diverse domains Zou et al. (2024). However, the sheer scale of LLM parameters renders traditional full-model fine-tuning prohibitively expensive. To address this challenge, various parameter-efficient fine-tuning (PEFT) techniques have been proposed, including prefix-tuning Li and Liang (2021), prompt-tuning Lester et al. (2021), and adapter-based methods Hu et al. (2023). Among these approaches, low-rank adaptation (LoRA) Hu et al. (2022), which leverages low-rank matrices to re-parameterize pre-trained weight matrices, has received unprecedented attention. LoRA requires tuning less than 1% of the parameters needed for a full fine-tune while still achieving comparable performance across a wide range of downstream tasks, without introducing additional inference latency. Building on these advantages, numerous LoRA variants have been developed to further improve its efficiency and accuracy Kopiczko et al. (2023); Zhang et al. (2023); Liu et al. (2024d).
@@ 2.2 Federated Fine-tuning of LLMs
Federated learning has attracted substantial research interest Li et al. (2025), serving as a paradigm for addressing data privacy concerns Liu et al. (2024b). Recently, federated fine-tuning of LLMs has gained significant attention, with most existing studies focusing on integrating PEFT methods into the FL framework to reduce computational costs Che et al. (2023); Wu et al. (2024); Cho et al. (2024); Babakniya et al. ; Zhang et al. (2024); Sun et al. (2024); Bai et al. (2024); Liu et al. (2024c). For example, Zhang et al. (2024) incorporates LoRA into the FedAvg framework so that only LoRA modules need to be trained and aggregated. Extending this approach to resource-constrained and heterogeneous scenarios, Wang et al. (2024) propose a stacking-based aggregation strategy for heterogeneous LoRA modules, where individual LoRA modules are uploaded for aggregation, and the resulting stacked full-size LoRA weights are distributed back to clients. Sun et al. (2024) further enhances performance under differential privacy guarantees and improves computational efficiency by fine-tuning only the zero-initialized LoRA matrices. Although these approaches reduce both computation and communication costs compared to full fine-tuning, transmitting LoRA modules still imposes considerable overhead. Even though LoRA accounts for a small portion of the total parameters, the massive scale of LLMs means these modules remain large. Repeatedly exchanging them during multiple rounds results in prohibitively high communication costs, making communication the dominant bottleneck in training time.
Another line of research leverages zeroth-order optimization methods for federated LLM fine-tuning Qin et al. (2024); Xu et al. (2024). While these approaches improve communication efficiency, their reliance on zeroth-order optimization significantly reduces computational efficiency compared to backpropagation-based methods. Consequently, these techniques substantially increase the computation time and prolong the overall training process, particularly in scenarios with limited clients or resource-constrained environments.
@@ 2.3 Communication Optimization in FL
Federated fine-tuning of LLMs, however, presents new challenges distinct from traditional FL, where PEFT techniques widely adopted in this context could lead to different training dynamics and parameter distributions. As a result, conventional optimization techniques, such as top-$k$ sparsification, may fail to exploit these unique properties, yielding suboptimal communication savings. Similarly, approaches like active client sampling Tang et al. (2022) often incur substantial computational overhead, which undermines their practicality in large-scale LLM fine-tuning scenarios.
@@ 3 Method
3.1 Problem Formulation
We consider an FL setting with one server and $K$ devices. Each device $i$ holds a local dataset $\mathcal{D}_{i}={(x_{j},y_{j})}^{n_{i}}$, where $n_{i}$, $x_{j}$, $y_{j}$ denote the number of samples, the input samples, and labels in client $i$, respectively. The total number of samples across all devices is $N=\sum_{i=1}^{K}n_{i}$. Following recent state-of-the-art approaches, the pre-trained LLMs $\mathcal{M}$ remain fixed on each device, while only the LoRA parameters are updated and exchanged between device $i$ and the server. Suppose $\mathcal{L}(\mathcal{M},\mathcal{P},x_{j},y_{j})$ is the loss evaluated by the model $\mathcal{M}$ with LoRA parameters $\mathcal{P}$ on the local data $(x_{j},y_{j})$, then the optimization goal is to find a set of LoRA parameters $\mathcal{P}$ to minimize the loss:
LoRA models the weight update $\Delta W\in\mathbb{R}^{m\times n}$ through a low-rank decomposition $BA$, where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are two low-rank matrices with $r\ll\min(m,n)$.
@@ 3.2 System Model
Our primary objective is to enhance communication efficiency specifically for federated fine-tuning of LLMs. To guide the design of such methods, we establish the following system goals:
@@
  • Robustness to Non-IID Data: Because data distributions can vary significantly across clients in real-world settings, our framework should remain robust under non-IID conditions.
To address these challenges, we propose EcoLoRA, a novel communication-efficient FL framework illustrated in Figure 1. First, we propose a round-robin segment-sharing scheme, leveraging the modular structure of LoRA. Instead of transmitting the entire LoRA module, each client shares only a complementary portion, significantly reducing communication overhead. Second, we introduce an adaptive sparsification tailored for the different training dynamics observed in matrices A and B. This method dynamically compresses parameters based on their training behavior, ensuring minimal performance degradation. Third, the adaptive sparsification method naturally enables a parameter distribution suitable for geometric compression, which we exploit through Golomb coding Golomb (1966) to further optimize communication efficiency. We elaborate round-robin segment sharing, adaptive sparsification, and encoding in Sections 3.3, 3.4, and 3.5, respectively. Additionally, we analyze computational overhead in Section 3.6 and provide convergence analysis in Section 3.7.
@@ 3.3 Round-Robin Segment Sharing
For example, consider $N_{t}=5$ clients and $N_{s}=3$ segments. In round $t=0$, client 0 uploads the segment with ID $(0+0)\bmod 3=0$, i.e., $s_{0,0}^{0}$; client $1$ uploads $s_{1,1}^{0}$; client $2$ uploads $s_{2,2}^{0}$; client $3$ uploads $s_{3,0}^{0}$; and client $4$ uploads $s_{4,1}^{0}$. The server then averages $s_{0,0}^{0}$ and $s_{3,0}^{0}$ to form the 0-th segment, averages $s_{1,1}^{0}$ and $s_{4,1}^{0}$ to form the 1-th segment, and takes $s_{2,2}^{0}$ for the 2-th segment. Because each client transmits only a single segment in each round, this round-robin segment sharing scheme reduces the upload communication load to $1/N_{s}$ of the total parameters.
However, this partial update approach introduces a delay for segments that are not uploaded in a given round, which can increase the number of rounds required to converge. To mitigate potential accuracy degradation, we leverage the local model by taking a weighted average of the global and local models at the beginning of each round before optimization. This ensures that even if a segment is not uploaded in a particular round, its previous state still guides local optimization. Moreover, by mixing the globally shared model (the consensus among clients) with the client’s locally fine-tuned model (adapted to its specific data), we improve robustness under non-IID distributions. In cross-device settings, only a subset of clients participates in each round, which may result in some clients remaining idle for many rounds and thus suffering from stale local parameters that potentially hamper global convergence Xie et al. (2019) when using the simple average. To address this, we employ an exponential decay weighting Chen et al. (2019) to update the local LoRA model:
@@ 3.4 Adaptive Sparsification
where $k^{t}$ is the sparsity level for round $t$, $L_{0}$ is the initial loss, $L_{t-1}$ is the global loss for round t1t-1, and $k_{\max}$ and $k_{\min}$ define the sparsification range. As training loss decreases, $k^{t}$ is reduced, reflecting that the model has learned sufficient knowledge and updates have become sparser. Second, to address the distinct patterns in matrices A and B, we introduce a matrix-adaptive sparsification scheme. We set smaller $k_{\min}$ value for B (due to its higher sparsity) and use a larger $\gamma$ for B to capture its rapid change in sparsity.
@@ 3.5 Lossless Encoding
[…]
@@ 4.1 Experimental Setup
FL Settings and Implementation Details. Following Zhang et al. (2024), we implement our framework in a federated learning environment with 100 clients. In each round, we randomly sample 10 clients and conduct training for 40 global rounds. To simulate realistic scenarios, we adopt a non-IID data distribution across clients. Detailed experimental configurations and hyperparameter settings are provided in Appendix A.
@@ 4.3 Evaluation in Practical Networks
To evaluate the performance of EcoLoRA under realistic network conditions, we implemented a simulated federated learning platform following Ekaireb et al. (2022), using ns-3, a widely adopted discrete-event simulator for network communications Henderson et al. (2008). Our simulation adopts ns-3’s point-to-point model to emulate realistic client–server TCP communication. We configure the TCP stack to match standard Linux implementations Sarolahti and Kuznetsov (2002).
Following practical uplink (UL) and downlink (DL) bandwidth settings in Konečnỳ (2016), we simulate four bandwidth scenarios: 0.2/1 Mbps, 1/5 Mbps, 2/10 Mbps, and 5/25 Mbps, with a fixed latency of 50ms to capture different network conditions. Figure 3 compares the computation and communication time of EcoLoRA against baselines under these scenarios, using Llama2-7B trained on Dolly. Our results demonstrate that as network conditions deteriorate, communication time increasingly dominates the total training time. This effect is particularly notable given that actual throughput typically falls short of theoretical bandwidth; for instance, in our simulated environment, a 1Mbps bandwidth connection achieved an average throughput of only 0.89Mbps. These findings underscore the importance of developing communication-efficient fine-tuning methods. Across all conditions, EcoLoRA significantly reduces communication overhead while introducing minimal computational cost. For instance, under the 1/5 Mbps setting, it reduces communication time by 79% and total training time by 65%. Moreover, the additional per-round computation cost remains below 3s, making EcoLoRA a practical solution for resource-constrained environments.
@@ 4.4 Ablation Study
In this section, we analyze the impact of various design components and hyperparameter choices. We present additional experiments in Appendix C.
@@
Threshold $k$ Fixed Top-$k$ Adaptive Sparsification
0.9 66.5 66.6
0.7 66.1 66.5
0.6 66.1 66.5
0.5 65.8 66.3
Number of Clients. We further examine the impact of scaling the total number of clients. To this end, we evaluate our method with two additional client populations using LLaMA2-7B fine-tuned on Alpaca. The results in Table 6 show that EcoLoRA consistently reduces communication costs while maintaining accuracy across different client scales.
Table 6: Comparison of accuracy and parameters (in millions) under different numbers of clients.
# Clients Method ARC Upload P. Total P.
200 FedIT 66.5 3858.8 7633.6
200 w/ EcoLoRA 66.5 501.9 2968.2
300 FedIT 66.3 4529.8 8933.9
300 w/ EcoLoRA 66.4 750.6 4450.0
Number of Clients Participating in Each Round. The number of clients participating in each communication round is another critical factor in FL. We fix the total number of clients to 100 and vary the number of participants per round. Experiments with LLaMA2-7B fine-tuned on Dolly (Table 7) show that EcoLoRA remains effective under different participation levels.
Table 7: Comparison of accuracy and parameters (in millions) under varying numbers of participating clients in each round.
# P. Clients Method ARC Upload P. Total P.
30 FedIT 66.5 4781.5 9437.2
30 w/ EcoLoRA 66.6 750.5 4449.9
50 FedIT 66.5 4613.7 9017.8
50 w/ EcoLoRA 66.5 645.9 3791.6
Number of Local Updates. We next study the impact of varying the number of local updates before communication. Experiments with LLaMA2-7B fine-tuned on Dolly (Table 8) show that EcoLoRA consistently achieves significant communication reduction while preserving accuracy across different local computation settings.
Table 8: Comparison of accuracy and parameters (in millions) under different numbers of local updates.
# Local Updates Method ARC Upload P. Total P.
3 FedIT 66.4 1342.2 2642.4
3 w/ EcoLoRA 66.4 151.0 880.8
5 FedIT 66.3 1006.6 1971.3
5 w/ EcoLoRA 66.4 110.7 639.2
Impacts of LoRA’s Rank. We also investigate the effect of LoRA rank using FFA-LoRA with two different ranks on LLaMA2-7B fine-tuned with Dolly. As shown in Table 9, EcoLoRA consistently delivers substantial communication savings at both low and high ranks, while maintaining accuracy.
Table 9: Comparison of accuracy and parameters (in millions) under different LoRA ranks.
LoRA Rank Method ARC Upload P. Total P.
8 FFA-LoRA 66.3 713.0 1405.1
8 w/ EcoLoRA 66.3 86.6 622.5
32 FFA-LoRA 66.7 2600.4 5117.0
32 w/ EcoLoRA 66.8 333.5 2567.5